[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Apr 20, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Apr 20 19:28:19 UTC 2017


HSU Shift Log for Thu Apr 20 2017 05:17:55 GMT-0400 (EDT)

Shift: A

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: Had to once again submit HWRF/GFDL CARQ twice at 06Z in order  
for SDM to receive it. No problems with guidance retrieval.

N-AWIPS issues: No GOES-R/-16 satellite data ingest in N-AWIPS since  
19/1700Z.

AWIPS-2/NCP issues: None. GOES-R/-16 satellite data ingest working properly.

Other Techincal issues: No ECMWF-based SHIPS available at 06Z despite  
pulling files 5 times.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170420/03d12881/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list